Eysian Fields, Inc, uses a maxmum payback period of 6 years and currenty must choose between two mutually exclusive projects. Project outsy of $29,000 project Helum requires an initial outlay of $35,000 Using t projects payback period Which project meets Elysiahs standards? ydrogen requires an initial the expected cash intows given for each project in the following table, EB. calculate each The payback pernod of project Hydrogen is years (Round to two decimal places) The payback penod of project Helium is years. (Round to two decimal places ) Which project meets Elysians standard? (Select the best answer below) O Only project Hydrogen meets Elysians standard O Only project Helium meets Elysians standard O Both projects are acceptable because their payback perniods are less than the 6 years criteriosn (Click on the icon located on the top-right comer of the data table below in order to copy its contents into a Year HydrogenHelium $7,000 $7 500 $3,500 $3,500 $2,500 $8,000 $7.500 $4.500 $6,000 $5,000 Clack to select your lype here to
Cost of debt using the approximation formula For the following $1,000 par-value bond, assuming annual int after-tax cost to maturity using the approximation formula Discount (-)or Coupon Life Underwriting fee premium () interest rate years $20 $50 7% The after-tax cost of fina ang using the approximation formulas % (Round toto decr al places Enter your answer in the answer box. 0
Answer #1

Payback period is the period after which the cumulative cash flows become positive

Project Hydrogen -

Year CF Cumulative
0 -29000 -29000
1 6500 -22500
2 7000 -15500
3 7500 -8000
4 3500 -4500
5 3500 -1000
6 2500 1500

The cash flow becomes positive after period 5

Hence, payback period = 5 + 1000/2500 = 5.4 years

Payback period for project Hydrogen = 5.4 years

Project Helium -

Year CF Cumulative
0 -35000 -35000
1 8000 -27000
2 8000 -19000
3 7500 -11500
4 4500 -7000
5 6000 -1000
6 5000 4000

The cash flow becomes positive after period 5

Hence, payback period = 5 + 1000/5000 = 5.2 years

Payback period for project Helium = 5.2 years

Since Elysian wants a payback period of less than 6 years, (c) Both projects are acceptable because their payback period are less than 6 years criterion